Expand and simplify each function. Evaluate at f(- 2).


a. f(x) = (x + 1)(x - 2)(x + 3)


b. f(x) = (2x ^ 2 - 3)(x ^ 3 + 1)


1
Expert's answer
2021-02-16T09:40:51-0500

(a) "f(x) = (x+1)(x-2)(x+3)"

"\rightarrow" "(x^2 -2x + x -2)(x+3)"

"\rightarrow (x^2 -x -2)(x+3)"

"\rightarrow x^3 + 3x^2 - x^2 -3x -2x -6"

"\rightarrow x^3 +2x^2 -5x -6"

For f(-2), put x= -2

"f(-2) = (-2)^3 +2(-2)^2 -5(-2) -6"

"\rightarrow -8 +8 +10 -6" = 4

So f(-2) = 4


(b) "f(x) = (2x^2 -3)(x^3 +1)"

"= (2x^5 +2x^2 -3x^3 -3)"

"=""2x^5 -3x^3 +2x^2 -3"


For f(-2), put x= -2

"f(-2) = 2(-2)^5 -3(-2)^3 +2(-2)^2 -3"

"= -64 +24 +8 -3"

"= -35"


So f(-2) = -35